ERRATUM TO “ON THE EXISTENCE OF PERIODIC SOLUTIONS

FOR A CLASS OF GENERALIZED FORCED LIÉNARD EQUATIONS” [APPL. MATH. LETT. 20 (2007), NO. 3, 248–254]

M. R. POURNAKI AND A. RAZANI

The authors would like to thank Professor Alessandro Fonda for pointing out the error in the statement of the Main Theorem. On pages 1 and 2 “T -periodic real function” should read “real function”. Also the statement of the Main Theorem should read as follows:

Main Theorem. Suppose f, k, and g are real functions on ℝ which are locally Lipschitz and p is a nonconstant, continuous, real function on [0, T], T > 0. Also suppose all solutions of the initial value problem (1.1) can be extended to [0, T]. If there exist real numbers a_{1} and a_{2} for which g(a_{1}) ≤ p(t) ≤ g(a_{2}) holds for each 0 ≤ t ≤ T, then there exists T_{0}, 0 < T_{0} < T, such that if p is T_{0}-periodic, Eq. (1.1) has at least one T_{0}-periodic solution.

M. R. Pournaki, Department of Mathematical Sciences, Sharif University of Technology, P.O. Box 11155-9415, Tehran, Iran, and School of Mathematics, Institute for Studies in Theoretical Physics and Mathematics, P.O. Box 19395-5746, Tehran, Iran.

  E-mail address: pournaki@ipm.ir
  URL: http://math.ipm.ac.ir/pournaki/

A. Razani, Department of Mathematics, Faculty of Science, Imam Khomeini International University, P.O. Box 34194-288, Qazvin, Iran, and School of Mathematics, Institute for Studies in Theoretical Physics and Mathematics, P.O. Box 19395-5746, Tehran, Iran.

  E-mail address: razani@ikiu.ac.ir
  URL: http://math.ipm.ac.ir/razani/